Week of 060529

Open Actions from last week:

Chair: James

Monday:

Log: LFC host certs ran out on saturday - half hosts down for 19 hours

New Actions:

  • look at excessive LOG_DB_ERRORs on LfC (Jamesc) DONE

Discussion:

Tuesday:

Log: nothing to report

New Actions:

Discussion:

  • Validation of LFC/FTS on 10gR2 done. Possible intervention tomorrow - waiting on final decision

Wednesday

Log:RAS

Actions:

Discussion:

Thursday

Log: RAS

Actions:

  • Olof mentioned that 15K files exist on SC4 pools, with no migrator turned on. Check if any SC related user created them (James)

Discussion:

Friday

Log: Problems with MDS on ce102 - crashed several times

Actions:

  • Investigate ce102 (Maarten)
  • Look at cron job error messages on new RB about proxycache directory (Maarten) DONE
  • Install new vomscert on LFC/FTS (SMOD)
  • Update GOCDB with new LFC endpoints (Harry) DONE
Discussion:

Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r3 - 2006-06-06 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback